OpenStudy (anonymous):

What is the approximate volume of a cone with a height of 3 meters and base radius of 8 meters? Use three point one four for pi. two hundred point nine six cubic meters two hundred twenty-six point zero eight cubic meters two hundred seventy-six point three two cubic meters six hundred two point eight eight cubic meters

OpenStudy (anonymous):

@RaeDeSol can you help me with this one also

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

V=πr^2(h/3)

OpenStudy (anonymous):

(3.14)(8^2)(3/3) = V (3.14)(64)(1) (200.96)

OpenStudy (anonymous):

TY

OpenStudy (anonymous):

You're welcome